# KTester: Leveraging Domain and Testing Knowledge for More Effective LLM-based Test Generation

### v1

@@ 3.2.1. Design
We use the HITS dataset (Section 3.1.1), which contains 110 test generation tasks from 10 open-source Java projects. Our method builds the project knowledge base by analyzing the corresponding project source code versions in the dataset. For these tasks, both KTester and baseline methods generate unit tests using GPT-4o-mini as the backbone model. For fair comparison, KTester and all baselines generate exactly one test class per focal method, with no restriction on the number of test cases or generation time within each class. This allows complex methods to achieve high coverage through diverse test cases. We limited automatic repair iterations to 5 in KTester. To mitigate the effect of randomness introduced by LLMs, we repeated each experiment three times and report the average results.
To systematically evaluate the effectiveness of generated unit tests, we adopt eight widely used metrics for unit test quality, capturing correctness, sufficiency and efficiency  (yuan2023no; wang_hits_2024; chen_chatunitest_2024; gu_testart_2025).
@@
Coverage is measured using Jacoco (Jacoco), an open-source tool for measuring code coverage in Java projects, ensuring objective and consistent evaluation. By combining correctness (CPR, EPR), coverage metrics (LC, BC, LCP, BCP) and efficiency metrics (AvT, AvTC), we provide a balanced and robust assessment of test quality, facilitating fair comparison between KTester and baseline approaches.
@@ 3.2.2. Results
Correctness. KTester reaches a perfect Compile Pass Rate (CPR) of 100%, indicating that all generated test classes are syntactically valid and type-correct. While UTGen also achieves 100% CPR, the other LLM-based baselines fall short, suggesting potential issues in code generation or dependency handling. Regarding the Execution Pass Rate (EPR), which measures runtime correctness, KTester attains 90.05%, slightly lower than UTGen. This is expected, as UTGen builds upon EvoSuite-generated tests, which provide a strong foundation for runtime correctness. Nevertheless, KTester still surpasses all other LLM-based methods in EPR, demonstrating its effectiveness in generating executable and correct test cases. This result indicates that incorporating project knowledge yields substantial benefits, as it more effectively guides LLMs toward generating executable test code.
Sufficiency. In terms of code coverage, KTester achieves the highest Line Coverage (LC) of 62.78% and Branch Coverage (BC) of 54.71%, significantly outperforming HITS (49.74% LC, 43.74% BC) while generating less test case (7.33 vs. 15.78). ChatTester yields the lowest coverage, highlighting its limited ability to explore program logic. UTGen is constrained by EvoSuite’s inability to operate on certain project-specific packages (e.g., com.apple.spark in the batch-processing-gateway project) and by the path-explosion problem inherent to SBST techniques. As a result, its coverage exceeds only that of ChatTester. When restricting evaluation to only the tests that both compile and execute successfully, KTester maintains the lead with 54.49% Line Coverage of Passed Tests (LCP) and 46.23% Branch Coverage of Passed Tests (BCP), confirming the practical adequacy of its outputs.
These results confirm that integrating project-specific and test-domain knowledge into LLM-based generation pipelines can significantly enhance the validity and adequacy of produced test cases. In particular, KTester demonstrates not only higher syntactic and semantic correctness, but also stronger capability in exercising program logic and control flow, making it a more effective and practical solution for automated unit test generation.
@@
Figure 10. Test Classes generated by HITS and KTester.
Generalizability. To assess whether KTester maintains its performance when implemented with alternative LLMs, we implemented KTester using claude-3.5-haiku-20241022 (claude-35-haiku) and deepseek-v3.1 (deepseek-v31), and the experimental results are reported in Table 3, which achieved even better results than gpt4o-mini, confirming its generalizability.
Finding 1: KTester outperforms all baselines across eight metrics. Compared to the strongest baseline (HITS), and while generating on average 8.45 fewer test cases, it improves execution pass rate by 5.69% and line coverage by 8.83%, demonstrating clear gains in both correctness and sufficiency.
@@
EPR 77.07 81.75 82.41
LC 61.10 66.46 71.22
BC 52.59 60.66 65.75
LCP 54.49 56.30 65.19
BCP 46.23 50.76 59.33
@@ 3.3.2. Results
Table 4 shows the performance of each KTester variant compared to the original method. All variants experience a drop across correctness and coverage metrics, confirming that each key component positively contributes to overall test quality. Among them, KTester-DGT, which removes the intermediate structured test case generation step, causes the most significant degradation. For example, Execution Pass Rate (EPR) drops by about 24.08% (from 77.07% to 52.99%) and Line Coverage (LC) decreases by approximately 12.61% (from 61.10% to 48.49%). This indicates that separating test case design from test method generation is crucial for effectiveness. The other variants—KTester-UTE, KTester-FMR, and KTester-MVG—also reduce performance, but their impacts are less distinct and harder to differentiate clearly. This suggests that while invocation pattern extraction, similar function retrieval, and multi-view prompting all improve results, the modular test case transformation step plays the most critical role.
Finding 2: The modular test case transformation (KTester-TCG) has the largest impact on performance, with EPR and LC dropping 22.15% and 14.29%, respectively. Other components also contribute positively but with less distinguishable effects.
@@
KTester-UTE 100 58.20 56.84 50.48 44.54 38.90
KTester-FMR 98.15 53.91 58.19 50.95 45.40 39.41
KTester-MVG 100 64.57 54.99 48.59 43.80 38.89
KTester-DGT 97.76 52.99 48.49 44.05 35.79 31.99
KTester 100 77.07 61.10 52.49 54.49 46.23
@@ 3.4.1. Design
Each question is rated on a four-point Likert scale—strongly agree, somewhat agree, somewhat disagree, and strongly disagree. This setup enables quantitative comparison of the subjective quality of tests generated by different approaches and allows us to assess whether our method consistently produces more developer-friendly outputs. Following prior work (yuan2023no; icse25PanKKPS25), we adopt an even-numbered scale to avoid ambiguous “neutral” responses; specifically, the four-point design clearly separates two positive options from two negative ones.
@@ 3.4.2. Results
Interestingly, although UTGen leverages a LLM to improve the readability of test code, it still received more than 67% negative responses across all three dimensions. We attribute this outcome to two main factors: (1) In many cases, some function signatures and variable names are not modified by the LLM, leaving the test code difficult for participants to understand. (2) EvoSuite invokes private methods through public caller methods, whereas LLM-based methods rely on reflection. When the source code changes, participants found it difficult to determine the scope of impact for the test cases produced by UTGen, which negatively affected their assessments of correctness and maintainability.
Finding 3: KTester consistently outperforms all baselines in human evaluation across correctness, readability, and maintainability.
@@ 4.1. LLM-based Unit Test Generation
Large language models (LLMs) have become a powerful tool for automated unit test generation, offering human-like test code that overcomes the readability and maintainability limitations of traditional methods (harman2015achievements; mcminn2011search; baldoni2018survey; cadar2008klee; cha2012unleashing). Existing LLM-based approaches can be broadly categorized into two paradigms: fine-tuning and prompt-based methods.
Fine-tuning approaches treat test generation as a supervised learning task, typically formulated as a sequence-to-sequence translation problem. A representative example is AthenaTest (tufano2020unit), which fine-tunes BART models on curated datasets where focal methods are inputs and complete test cases are outputs. These models learn explicit mappings between source code and corresponding tests using domain-specific corpora. Variants of this paradigm (dakhel2024effective; deng2024large; deng2023large), including TeCo (nie2023learning), explore different encoder-decoder architectures to better capture the semantic relationships between code and the associated test logic. A recent example is EXLONG (zhang2024exlong), fine-tuned from CodeLlama to generate exceptional behavior tests. Unlike such fine-tuning approaches, our method injects project and testing knowledge via prompt design and generation workflow, making it more flexible and adaptable to stronger future LLMs.
Prompt-based methods, in contrast, rely on the zero-shot reasoning abilities of instruction-tuned models without additional training. Systems such as ChatTester (yuan2023no), ChatUniTest  (chen_chatunitest_2024) and ASTER (icse25PanKKPS25) use carefully designed prompts to guide models like GPT in generating test suites. These methods often adopt a multi-step reasoning process, decomposing tasks into subcomponents such as test identification, input generation, and assertion formulation. TestPilot (schafer2023empirical) enhances this strategy by incorporating documentation and usage examples, while TestSpark (sapozhnikov2024testspark) combines prompt-based generation with search-based techniques in IDEs.
Recent work explores hybrid approaches that integrate multiple strategies. CODAMOSA (lemieux2023codamosa) combines LLM-generated test seeds with evolutionary algorithms to improve coverage-driven test generation. CoverUp (altmayer2025coverup) employs iterative dialogues to refine outputs based on coverage feedback. SymPrompt  (ryan2024code) leverages symbolic execution to guide prompting.
Despite these advancements, major challenges remain. Most methods lack access to project-specific context, such as class instantiation, module dependencies, and API usage patterns—leading to syntactically correct but semantically invalid tests. Moreover, they often ignore foundational testing principles, including boundary value analysis, equivalence partitioning, and exception handling, resulting in tests that may execute without error but fail to thoroughly validate functionality.
@@ 4.2. Knowledge-enhanced LLMs for SE Tasks
Recent studies have explored enhancing LLM performance on SE tasks by incorporating task-relevant context. In APR, Xia et al. (Xia2023) showed that simply providing the buggy function can outperform traditional tools, while ChatRepair (XiaAuto2024) leveraged failing test names and assertions for interactive prompting. Further work has demonstrated the effectiveness of enriching context with bug-localized code (Prenner2024), relevant identifiers (XiaPlastic2024; JiangImpact2023), stack traces (haque-etal-2025-towards), and bug reports (FakhouryNL2Fix2024).
In parallel, RAG has emerged as a promising paradigm for integrating external knowledge into LLMs by retrieving relevant information from codebases or databases. For code completion, ReACC (lu2022reaccretrievalaugmentedcodecompletion) retrieves semantically similar code snippets, and Wu et al. (WuREPOFORMER2024) propose a selective RAG framework to reduce redundant retrievals. In code generation, recent approaches (li2025coderagsupportivecoderetrieval; Athale_2025) model code repositories as graph-based or knowledge-structured representations to retrieve and incorporate repository-level context into LLMs.
@@ 5. Threats to Validity
A key threat lies in the use of LLMs. To ensure fairness, all approaches are evaluated using the same model version. For baselines, we rely on official implementations or carefully reimplement them following published details, verifying outputs to align with reported results. Benchmark tasks and ground-truth tests are directly reused from prior work to ensure quality and consistency.
Another threat concerns generality and subjectivity. Our experiments focus on Java to remain consistent with prior work and enable direct comparison (yuan2023no; wang_hits_2024), but KTester is not inherently tied to Java. Extending to other frameworks mainly requires updating prompts and replacing libraries, while supporting other languages involves only substituting the AST analysis tool (e.g., tree-sitter). Thus, the design is modular, and the adaptation cost is low. Future work will explore such extensions. While correctness and coverage provide objective metrics, they may not fully capture clarity and structure. We mitigate this via a user study on readability and maintainability, though results may still reflect participant bias. The study includes 14 professional developers, which we believe is sufficient, but broader sampling could further strengthen validity.
@@ 6. Conclusions
In this paper, we present KTester, a project-aware and testing-domain-knowledge-guided framework for LLM-based unit test generation. By extracting comprehensive project knowledge and decoupling test case design from test code generation, KTester effectively guides the LLM to produce semantically rich and maintainable test cases. Our multi-perspective prompting strategy ensures diverse and thorough coverage of testing scenarios, while structured test templates enhance code clarity and reusability. Extensive evaluations on real-world open-source projects demonstrate that KTester consistently outperforms existing state-of-the-art methods in correctness, coverage, readability, and maintainability. Future work includes exploring dynamic analysis to further enrich project knowledge and extending the approach to other programming languages and testing paradigms.

### v2

@@ 3.2.1. Design
We use the HITS dataset (Section 3.1.1), which contains 110 test generation tasks from 10 open-source Java projects. Our method builds the project knowledge base by analyzing the corresponding project source code versions in the dataset. For these tasks, both KTester and baseline methods generate unit tests using GPT-4o-mini as the backbone model. For fair comparison, KTester and all baselines generate exactly one test class per focal method, with no restriction on the number of test cases or generation time within each class. This allows complex methods to achieve high coverage through diverse test cases. We limited automatic repair iterations to 5 in KTester. To mitigate the effect of randomness introduced by LLMs, we repeated each experiment five times and report the average results.
To systematically evaluate the effectiveness of generated unit tests, we adopt eight widely used metrics for unit test quality, capturing correctness, sufficiency and efficiency  (Yuan et al., 2024; Wang et al., 2024; Chen et al., 2024; Gu et al., 2025).
@@
Coverage is measured using Jacoco (33), an open-source tool for measuring code coverage in Java projects, ensuring objective and consistent evaluation. By combining correctness (CPR, EPR), coverage metrics (LC, BC, LCP, BCP) and efficiency metrics (AvT, AvTC), we provide a balanced and robust assessment of test quality, facilitating fair comparison between KTester and baseline approaches.
@@ 3.2.2. Results
Correctness. KTester reaches a perfect Compile Pass Rate (CPR) of 100%, indicating that all generated test classes are syntactically valid and type-correct. While UTGen also achieves 100% CPR, the other LLM-based baselines fall short, suggesting potential issues in code generation or dependency handling. Regarding the Execution Pass Rate (EPR), which measures runtime correctness, KTester attains 77.07%, slightly lower than UTGen. This is expected, as UTGen builds upon EvoSuite-generated tests, which provide a strong foundation for runtime correctness. Nevertheless, KTester still surpasses all other LLM-based methods in EPR, demonstrating its effectiveness in generating executable and correct test cases. This result indicates that incorporating project knowledge yields substantial benefits, as it more effectively guides LLMs toward generating executable test code.
Sufficiency. In terms of code coverage, KTester achieves the highest Line Coverage (LC) of 63.94% and Branch Coverage (BC) of 55.46%, significantly outperforming HITS (52.27% LC, 45.93% BC) while generating less test case (7.33 vs. 15.78). ChatTester yields the lowest coverage, highlighting its limited ability to explore program logic. UTGen is constrained by EvoSuite’s inability to operate on certain project-specific packages (e.g., com.apple.spark in the batch-processing-gateway project) and by the path-explosion problem inherent to SBST techniques. As a result, its coverage exceeds only that of ChatTester. When restricting evaluation to only the tests that both compile and execute successfully, KTester maintains the lead with 55.52% Line Coverage of Passed Tests (LCP) and 47.21% Branch Coverage of Passed Tests (BCP), confirming the practical adequacy of its outputs.
@@
EPR 76.41 81.75 82.41
LC 63.94 66.46 71.22
BC 55.46 60.66 65.75
LCP 55.52 56.30 65.19
BCP 47.21 50.76 59.33
Figure 10. Test Classes generated by HITS and KTester.
Test Classes generated by HITS and KTester.
These results provide evidence that integrating project-specific and test-domain knowledge into LLM-based generation pipelines is associated with improved test adequacy and correctness, as reflected by higher coverage and pass rates. In particular, KTester demonstrates stronger capability in exercising program logic and control flow, making it a more effective and practical solution for automated unit test generation.
Generalizability. To assess whether KTester maintains its performance when implemented with alternative LLMs, we implemented KTester using claude-3.5-haiku-20241022 (11) and deepseek-v3.1 (16), and the experimental results are reported in Table 3, which achieved even better results than gpt4o-mini, confirming its generalizability.
Finding 1: KTester outperforms all baselines across eight metrics. Compared to the strongest baseline (HITS), and while generating on average 8.45 fewer test cases, it improves execution pass rate by 5.03% and line coverage by 11.67%, demonstrating clear gains in both correctness and sufficiency.
@@ 3.3.2. Results
Table 4 summarizes the performance of KTester variants. All variants exhibit performance drops across correctness and coverage metrics, confirming that each component contributes to KTester. Notably, KTester-UTE shows the largest correctness degradation, with CPR and EPR decreasing by 6.37% and 14.41%, respectively, highlighting the critical role of usage trace extraction in improving test correctness. KTester-DGT leads to the most substantial coverage losses, with LC and BC dropping by 13.39% and 11.15%, respectively, indicating that separating test case design from test method generation is crucial for test adequacy. The other variants—KTester-FMR and KTester-MVG have comparatively smaller impacts across both correctness and coverage metrics, but also result in performance reductions.
Finding 2: Usage trace extraction (KTester-UTE) is the primary contributor to test correctness, while modular test case transformation (KTester-DGT) contributes most to test adequacy. Other components also contribute positively with smaller and less distinguishable effects.
@@
KTester-UTE 93.63 62.00 57.53 51.27 45.37 39.83
KTester-FMR 95.66 66.99 59.17 52.32 46.83 41.02
KTester-MVG 97.45 75.31 53.67 46.26 45.48 39.36
KTester-DGT 96.14 64.96 50.55 44.31 41.07 35.65
KTester 100 76.41 63.94 55.46 55.52 47.21
@@ 3.4.1. Design
Each question is rated on a four-point Likert scale—strongly agree, somewhat agree, somewhat disagree, and strongly disagree. This setup enables quantitative comparison of subjective test quality across approaches. Following prior work (Yuan et al., 2024; Pan et al., 2025), we adopt an even-numbered scale to avoid ambiguous “neutral” responses, and clearly separates positive from negative judgments.
Score Distribution of Correctness, Readability and Maintainability
@@ 3.4.2. Results
Interestingly, although UTGen uses an LLM to improve test readability, it still received more than 67% negative responses across all dimensions. This is mainly due to (1) function signatures and variable names often remaining unmodified, leaving the test code hard to understand, and (2) UTGen’s EvoSuite component testing private methods via public callers, while LLM-based methods rely on reflection, making participants feel hard to identify impacted tests after source code changes and thus negatively affecting assessments of correctness and maintainability.
Finding 3: KTester outperforms all baselines in human evaluation across correctness, readability, and maintainability.
@@ 4.1. LLM-based Unit Test Generation
Large language models (LLMs) have become a powerful tool for automated unit test generation, offering human-like test code that overcomes the readability and maintainability limitations of traditional methods (Harman et al., 2015; McMinn, 2011; Baldoni et al., 2018; Cadar et al., 2008; Cha et al., 2012). Existing LLM-based approaches can be broadly categorized into two paradigms: fine-tuning and prompt-based methods.
Fine-tuning approaches frame test generation as a supervised sequence-to-sequence task, learning mappings from focal methods to complete test cases using curated corpora, e.g., AthenaTest (Tufano et al., 2020), TeCo (Nie et al., 2023), and subsequent variants (Dakhel et al., 2024; Deng et al., 2024, 2023), which explore different encoder–decoder architectures to capture the semantic relationships between code and tests. EXLONG (Zhang et al., 2025) further fine-tunes CodeLlama for exception-oriented test generation. Unlike these approaches, our method injects project and testing knowledge through prompt design and generation workflows, enabling greater flexibility and compatibility with future LLMs.
Prompt-based methods, in contrast, exploit the zero-shot reasoning abilities of instruction-tuned models. Systems such as ChatTester (Yuan et al., 2024), ChatUniTest (Chen et al., 2024), and ASTER (Pan et al., 2025) use carefully designed prompts and multi-step reasoning to guide models like GPT in test identification, input generation, and assertion construction. Extensions incorporate additional context or search strategies, including documentation and usage examples in TestPilot (Schäfer et al., 2023) and IDE-based search in TestSpark (Sapozhnikov et al., 2024).
Recent work explores hybrid approaches that combine LLMs with complementary techniques. CODAMOSA (Lemieux et al., 2023) integrates LLM-generated test seeds with evolutionary algorithms to improve coverage-driven test generation, CoverUp (Altmayer Pizzorno and Berger, 2025) refines tests through coverage-guided dialogues, and SymPrompt (Ryan et al., 2024) leverages symbolic execution to steer prompting.
Despite these advancements, most approaches lack project-specific context (e.g., instantiation patterns, dependencies, and API usage) and overlook core testing principles such as boundary analysis, equivalence partitioning, and exception handling, yielding tests without execution error but insufficiently validate functionality.
@@ 4.2. Knowledge-enhanced LLMs for SE Tasks
Recent studies have explored enhancing LLM performance on SE tasks by incorporating task-relevant context. In APR, Xia et al. (Xia et al., 2023) show that simply providing the buggy function can outperform traditional tools, while ChatRepair (Xia and Zhang, 2024) leveraged failing test names and assertions for interactive prompting. Other work enriches context with bug-localized code (Prenner and Robbes, 2024), relevant identifiers (Xia et al., 2024; Jiang et al., 2023), stack traces (Haque et al., 2025), and bug reports (Fakhoury et al., 2024).
In parallel, retrieval-augmented generation (RAG) integrates external knowledge by retrieving relevant information from codebases or databases. For code completion, ReACC (Lu et al., 2022) retrieves semantically similar code snippets, and Wu et al. (Wu et al., 2024) propose a selective RAG framework to reduce redundancy. For code generation, recent work (Li et al., 2025; Athale and Vaddina, 2025) represents code repositories as graphs or structured knowledge to retrieve and add repository-level context to LLMs.
@@ 5. Threats to Validity
A key threat lies in the use of LLMs. To ensure fairness, all approaches are evaluated using the same model version. For baselines, we rely on official implementations or carefully reimplement them following published details, verifying outputs to align with reported results. Benchmark tasks and ground-truth tests are directly reused from prior work for consistency.
Another threat concerns generality and subjectivity. Although experiments focus on Java for comparability (Yuan et al., 2024; Wang et al., 2024), KTester is not inherently tied to Java. Extending to other frameworks mainly requires prompt and library updates, while supporting other languages only involves replacing the AST analysis tool (e.g., tree-sitter), resulting in low adaptation cost. Future work will explore such extensions. While correctness and coverage are objective metrics, they may not fully capture clarity and structure. we mitigate this via a user study on readability and maintainability with 14 professional developers, though results may still reflect participant bias and limited sampling. We believe this is sufficient, but a broader sampling could further strengthen validity.
@@ 6. Conclusions
We present KTester, a project-aware, testing-domain-knowledge-guided framework for LLM-based unit test generation. By extracting comprehensive project knowledge and decoupling test case design from test code generation, KTester effectively guides the LLM to produce semantically rich and maintainable tests. Our multi-perspective prompting strategy ensures diverse and thorough coverage of testing scenarios, while structured test templates enhance code clarity and reusability. Extensive evaluations on real-world open-source projects show that KTester consistently outperforms existing state-of-the-art methods in correctness, coverage, readability, and maintainability. Future work explores integrating dynamic analysis for enriching project knowledge and extending the approach to other languages and testing paradigms.
Acknowledgements.
This work was supported by National Natural Science Foundation of China (Grant No. 62402113), the General Program of the Natural Science Foundation of Guangdong Province, China (Grant No. 2025A1515011631), Social Science Planning Project (Grant No. SZ2025A002), and GMCC-SYSU Joint Lab for Smart Applications.